• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Carlos

    Log entry time 08:15:04 on December 1,2004

    Entry number 137589

    keyword=Owl shift summary



    Shift crew:
    ===========
    C. Munoz Camacho
    B. Moffit

    Shift summary:
    ==============

    00:00 Shift started with the calorimeter elastic calibration going on.

    02:00 Elastic calibration finished. Asked for controlled access to push calorimeter foward to running position. LHRS magnets cycled and set to 2 elastic kinematics. LHRS moved to the right angle.

    02:36 Elastic run at 1.1m on LH2.

    03:55 Elastic run on multifoil target, 1.1m. 10uA.

    04:00 Compton giving problems (DAQ crashing) at 10uA. We are basically not running it...

    06:45 Moved to carbon target. Elastic run at 10uA and 1.1m.

    07:08 Asked for 20uA. Rates were 3Hz and calorimeter ADCs quite low.

    07:35 Asked for controlled access to install veto.


    Run summary:
    ============
    Run 4994 : Elastic calibration run setting 1. PMTs warming up, 46k events
    Run 4995 : Elastic calibration run setting 1. 70k events
    Run 4996 : Elastic calibration run setting 2. 72k events
    Run 4997 : Elastic calibration run setting 3. 71k events

    Run 4998 : Elastic run at 1.1m. Setting 2. 50kevents. 2uA.
    Run 4999 : Elastic run at 1.1m on multifoil target. Setting 2. 10uA.
    Run 5000 : Elastic run at 1.1m on C target. Setting 2.


    Trip summary:
    =============
    00:00-01:00 19 trips (6min BNA)
    01:00-02:00 14 trips (5min BNA)
    02:00-03:00 14 trips (5min BNA)
    03:00-04:00 19 trips (6min BNA)
    04:00-05:00 20 trips (7min BNA)
    05:00-06:00 24 trips (8min BNA)
    06:00-07:00 26 trips (9min BNA)
    07:00-08:00 26 trips (9min BNA)